Write a system of two equations in two solutions for each problem. Solve each system by substitution.
Jay Lenos's garage. Jay Leno's collection of cars and motorcycles totals 187. When he checks the air in the tires he has 588 tires to check. How many cars and how many motorcycles does he own? Assume that the cars all have four tires and the motorcyles have two.

Answer by stanbon(75887)

Jay Lenos's garage. Jay Leno's collection of cars and motorcycles totals 187. When he checks the air in the tires he has 588 tires to check. How many cars and how many motorcycles does he own? Assume that the cars all have four tires and the motorcyles have two.
---------------------------
Cars Equation: c + m = 187
Tires Equation:4c + 2m = 588
---------------------------------
Multiply thru 1st by 4 to get:
4c + 4m = 4*187
---
Subtract the Tires Equation from that and solve for "m":
2m = 160
m = 80 (# of motorcycles)
---
Since c+m = 187, c = 107 (# of cars)
=======================================
cheers,
Stan H.
=======================================